Session: Feb 2017 (session 1)
Wavelength: 21 cm


To give feedback on this experiment, click here

Experiment Feedback

Badary (Plotnikova, Zhukova at Fri Feb 24 19:49:24 2017) Success:

Medicina (A.Mattana at Fri Feb 24 23:47:31 2017) Success:
No failures but Windy and Rainy Night

Zelenchukskaya (Lysenkova M. at Sat Feb 25 00:13:09 2017) Success:
Scans from no0052 to no0085 missed due to participation in intensive geo session

Urumqi (Hua Zhang at Sat Feb 25 08:36:49 2017) Minor failures:
18£º00£šUT£©,Hydrogen maser failure, replace the hydrogen maser.

Svetloe (Gusev D. at Sat Feb 25 08:43:07 2017) Success:

Westerbork (Blaauw at Sun Feb 26 08:54:36 2017) Success:

Onsala (Jun Yang at Mon Feb 27 09:17:18 2017) Success:
No known problems.

Effelsberg (Uwe at Wed Mar 01 13:14:24 2017) Success:
No known problems.

Robledo (Cristina Garcia Miro at Wed Mar 08 08:25:24 2017) Success:
During setup time found a problem with VEX file (prepared for year 2016), retrieved from vlbeer corrected VEX file, no impact. Recorded on Mark5 module MPI-0528, please note that VDIF files with name in capital letters without useful data, correct VDIF files in lower case.

Torun (Pawel Wolak at Fri Mar 17 11:06:27 2017) Success: HIGH

Jodrell_Bank (Jimmy Cullen at Thu Apr 27 12:59:17 2017) Success:
No known problems